Solve for x
x^2-12x+35=15

Answer by stanbon(75887)

Solve for x
x^2-12x+35=15
-----
x^2 -12x + 20 = 0
x^2-10x-2x+20 = 0
x(x-10)-2(x-10) = 0
(x-10)(x-2) = 0
x = 10 or x = 2
=====================
Cheers,
Stan H.
================